How much energy in kilo joule is produced when 7.1 gram of gaseous chlorine atoms are converted gaseous chloride ions?(electron affinity of chlorine is -3.7 eV.1eV/atm =96.49 kJ/moles)

Grade:10

1 Answers

Arun
6 years ago

Electron affinity in eV = 3.7

Electron affinity in kcal/ mol = 3.7 * 23.06

= 85.322 kcal / mol

Number of moles = 7.1 / 35.5 = .2

1 mole gives 85.322 kcal

0.2 gives 85.322 * 0.2

= 17.0644 kcal

= 17.0644* 4.2 k Joule

= 71.652 KJoule